OpenStudy (anonymous):

how do you know if 2^3 is 2x2x2=8 or 2x2x2=256

OpenStudy (anonymous):

just multiply it out 2^3 is 8 2 times 2 = 4 4 times 2 = 8

OpenStudy (anonymous):

3^3 = ??

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

An integer exponent refers to the number of factors of the base. In the equation below, b is the base and n is a positive exponent. Example 1 \[3^4=3*3*3*3\]note four factor of three. Example2\[2^5=2*2*2*2*2\]again five factor of the base two